OpenStudy (anonymous):

what would be the slope and y-intercept of the line 4x-y+6=0?

OpenStudy (anonymous):

separate y and see what you get

OpenStudy (anonymous):

so would it just be the equation without the y? or what x is equal to?

OpenStudy (anonymous):

no no i mean y= ???

OpenStudy (anonymous):

how would you find what y is equal to though?

OpenStudy (anonymous):

y=6

OpenStudy (anonymous):

ok see 4x-y+6=0 subtract 6 from both sides 4x-y = -6 subtract 4x from both sides -y=-4x-6 => y=4x+6 compare it with this equation y = mx +b where m = slope and b = y-intercept

OpenStudy (anonymous):

can you find slope and y intercept form this equation now??? y=4x+6

OpenStudy (anonymous):

slope equals 4x and y equals 6 then?

OpenStudy (anonymous):

not 4x , slope = 4 and y-intercept = 6

OpenStudy (anonymous):

oh ok, thx again btw